OpenStudy (anonymous):

HELP! 72 students in school can ski, this is 24% of enrollment. how many are enrolled? How do I do this

OpenStudy (mathteacher1729):

The big idea with percentage problems like these is "X percent means X over 100". \[32% = \frac{32}{100}\] \[5% = \frac{5}{100}\] etc... So now the problem says 72 students out of (???) is the same as 24 out of 100. So we set up this equation: \[\frac{72}{X}=\frac{24}{100}\] "72 out of (???) is the same as 24 out of 100." Now just cross multiply and solve for X. Hope this helps. :)

OpenStudy (mathteacher1729):

Oh dear, something went wrong with the math formatting... Here were the examples I was trying to type up: \[32\% = \frac{32}{100}\] \[5\% = \frac{5}{100}\] Hope this works better..